# Lattice simulations of axion-U(1) inflation: gravitational waves, magnetic fields, and black holes

### v1

@@ Abstract
We numerically study axion-U(1) inflation, focusing on the regime where the coupling between axions and gauge fields results in significant backreaction from the amplified gauge fields during inflation. These amplified gauge fields not only generate high-frequency gravitational waves (GWs) but also induce spatial inhomogeneities in the axion field, which can lead to the formation of primordial black holes (PBHs). Both GWs and PBHs serve as key probes for constraining the coupling strength between the axion and gauge fields. We find that, when backreaction is important during inflation, the constraints on the coupling strength due to GW overproduction are relaxed compared to previous studies, in which backreaction matters only after inflation. For PBH formation, understanding the probability density function (PDF) of axion field fluctuations is crucial. While earlier analytical studies assumed that these fluctuations followed a $\chi^{2}$-distribution, our results suggest that the PDF tends toward a Gaussian distribution in cases where gauge field backreaction is important, regardless whether during or after inflation. We also calculate the spectrum of the produced magnetic fields in this model and find that their strength is compatible with the observed lower limits.
@@ 1 Introduction
The idea of inflation originated from attempts to solve the singularity problem [1] and to address the horizon, isotropy and flatness problems in the hot big bang along with the monopole problem of grand unification, see [2, 3, 4] and [5, 6]. Inflation not only ameliorates the latter set of issues, but most importantly provides a natural mechanism [7, 8, 9] for generating initial density perturbations in the universe from mere vacuum fluctuations, see also [10, 11, 12, 13]. These perturbations account for anisotropies in the cosmic microwave background radiation (CMB) recently measured by Planck collaboration [14] with unprecedented precision, thereby confirming the predictions of this mechanism and the theory of inflation. Later on these perturbations collapse to form all galaxies and the large-scale structure in the universe. Following the initial proposal, numerous models have been suggested to realize an inflationary era in the early universe (for details see refs. [15, 16, 17, 18, 19]). Axion inflation models are particularly interesting because they provide a natural high energy physics ingredient for a long lasting inflation - an approximate shift symmetry of the potential [20]. After the first proposal, several models have been suggested in the context of axion inflation [21, 22, 23, 24, 25, 26]; see ref. [27] for a review. One of these models [25] considers a coupling of the axion field with the U(1) gauge field and we focus on this model in this paper. The axion inflation model, where the axion couples to a gauge field, exhibits a rich phenomenology. This includes the production of gravitational waves (GWs) [28, 29, 30, 31, 32, 33, 34], primordial black holes (PBHs) [35, 36, 37], matter-antimatter asymmetry [38, 39], and primordial magnetic fields [40, 41, 42, 43, 44, 45, 46, 47]. The rich phenomenology of axion inflation has prompted extensive studies to understand the dynamics of these models. In scenarios with a sufficiently strong coupling between the axion and the gauge field, the backreaction of the produced gauge fields becomes important either during inflation or post inflation depending on the coupling strength. This model including the backreaction of the gauge fields has been studied semi-analytically through perturbative approaches [48, 49, 50] and the gradient expansion formalism (GEF) [51, 47] under the assumption of a homogeneous axion field. More recently, efforts have been made to numerically investigate the strong backreaction regime, where the assumption of homogeneous axion field is relaxed [52, 53, 54]. Motivated by the importance of inhomogeneities of the axion field, as demonstrated in the numerical studies, these inhomogeneities have also been incorporated perturbatively into GEF [55].
@@
This rich phenomenology of axion inflation has been used to constrain the coupling strength between the axion and gauge fields. The coupling between the axion and the gauge field results in the production of gauge fields, which in turn generate high-frequency GWs, typically in the GHz range [33, 34]. These GWs contribute to the early universe’s total energy budget as additional radiation degrees of freedom. The bound on additional radiation degrees of freedom from CMB observations constrains the produced GWs, which further limits the coupling between the axion and the gauge fields. We reexamine the bounds on the coupling strength between the axion and the gauge field, considering stronger couplings than those explored in previous works [33, 34]. In [33], the authors considered scenarios where the backreaction due to the produced gauge field became important in the post-inflationary era but not during inflation. For one of their models, the chaotic inflation model, they found that the coupling strength should be below $(m_{\mathrm{pl}}/70)^{-1}$ for an inflaton mass equal [^1] Here and throughout the paper $m_{\mathrm{pl}}=1.22\times 10^{19}\,\,\rm{GeV}$ represents the Planck mass. to $1.2\times 10^{-6}m_{\mathrm{pl}}$. In our study, we explore scenarios with larger coupling strengths, where the gauge field energy density becomes large enough to affect the duration of inflation. Corresponding studies in refs. [52, 53] did not address the production of GWs, however. In these scenarios, inflation ends when the gauge field energy density becomes comparable to the inflaton field’s potential energy. Our findings indicate that in such scenarios, the bounds on the coupling strength are relaxed compared to previous estimates [33, 34].
@@
Additionally, the produced gauge field can also enhance spatial fluctuations of the axion field. If these spatial fluctuations exceed a threshold value, they can lead to the formation of PBHs222 For other models of the formation of PBHs during inflation see, e.g., refs. [56, 57, 58, 59]. in the post-inflationary era. This phenomenon has been analyzed in detail considering the $\chi^{2}$-distribution of the spatial fluctuations of the axion field. Given that the backreaction of the gauge field on the axion field is quadratic in nature, and assuming the gauge field has a Gaussian distribution, the spatial fluctuations of the axion field are expected to follow a $\chi^{2}$-distribution. By using constraints on PBH abundances, a bound on the axion and gauge field coupling has been obtained [35, 36]. A recent numerical study of axion inflation [52] considers the case when the backreaction of the gauge field on the axion evolution becomes significant during inflation. It was found that the distribution of the spatial fluctuations of the axion field starts to deviate from the $\chi^{2}$-distribution and begins to resemble a Gaussian distribution. In this paper, we study the probability distribution of spatial fluctuations in the axion field over a range of coupling strengths. We find that the distribution tends to a $\chi^{2}$-distribution during the evolution. However, it gradually transitions to a Gaussian distribution at a later time, irrespective of whether the backreaction of the gauge fields becomes important during or after inflation.
@@ 2 Axion-U(1) inflation model
Here, $\phi$ represents the axion field and $V(\phi)$ denotes its potential, which in our analysis is taken as [^3] Although chaotic inflation has been ruled out by CMB observations [18], we have chosen this model for several reasons. First, it simplifies our analysis. Second, it allows us to compare our results directly with those of previous studies. Lastly, many other inflationary potentials can be approximated by a quadratic potential near the end of inflation, and our study primarily focuses on this late phase of the inflationary period. $V(\phi)=m^{2}\phi^{2}/2$ with $m=1.06\times 10^{-6}m_{\rm pl}$, $F_{\mu\nu}=\partial_{\mu}A_{\nu}-\partial_{\nu}A_{\mu}$ represents the gauge field strength tensor, while $\tilde{F}^{\mu\nu}$ is its Hodge dual, $\alpha/f$ denotes the coupling strength between the axion field and the U(1) gauge fields, and $R$ is the Ricci scalar. Using this, and denoting derivatives with respect to conformal time $\eta$ by a prime, we get the following dynamical equations [^4] In this paper, we neglect the role of metric perturbations. The terms that couple metric perturbations to the axion field are suppressed by slow-roll parameters and can thus be ignored during inflation [60, 61]. However, the role of these terms may become important near the end of inflation, which we leave for future investigation; see refs. [62, 63, 54] for recent studies on this topic. for the axion-gauge field evolution [43]
@@ 3.1 Initial conditions
In this section, we describe the initial conditions for the axion and gauge fields used in our simulations. The initial conditions are set well within the inflationary era. To achieve accelerated expansion of the universe, the pressure $p$ must satisfy $p<-\rho/3$ which in the absence of gauge field implies $\phi^{\prime 2}<a^{2}V(\phi)$. Using this condition in its extreme form of $\ll$ in equations (2.5), we get
@@
here $H=\dot{a}/a=\mathcal{H}/a$ is the usual Hubble parameter, dot denotes derivative with respect to cosmological time $dt=ad\eta$ and $\epsilon_{H}$ represents the first slow-roll parameter. The above expressions imply
@@ 3.2 Dynamics of axion and gauge fields: inferred from simulations
To study the dynamics of the axion-U(1) inflation model, we perform a series of simulations with different coupling strengths [^6] It is to be noted that our coupling strengths are in the units of the inverse of the Planck mass. However, many previous studies use the inverse of the reduced Planck mass for normalization. To convert our coupling strengths to the reduced Planck mass normalization, our values need to be divided by $\sqrt{8\pi}$. table 1 provides the values of the coupling strengths in both units for clarity. $\alpha/f=30$, $50$, $65$, $75$, and $90$. The parameters for these simulations are given in table 1. In figure 1, we present the evolution of different energy densities for $\alpha/f=60$, $75$, and $90$. We show the evolution of the kinetic energy density, $\rho_{\mathrm{kin}}$ (dashed blue curves), potential energy density, $V(\phi)$ (solid cyan curves), and gauge field energy density, $\rho_{g}$ (dot-dashed orange curves) in panels (a), (b), and (c) of figure 1 for the coupling strengths $\alpha/f=60$, $75$, and $90$, respectively. Panels (d), (e), and (f) illustrate the evolution of $\xi$, while panels (g), (h), and (i) display the first slow-roll parameter, $\epsilon_{H}$, for the corresponding cases. Inflation ends when $\epsilon_{H}=1$. The vertical gray lines indicate the epoch of the end of inflation in each case. In this figure, $e$-folds $N$ is defined [^7] As usual we define dimensionless $e$-folds $N$ through the relation $dN=Hdt$ with cosmological time $t$ and Hubble parameter $H$. such that $N=0$ marks the end of inflation in scenarios where the backreaction of the gauge fields on the axion dynamics is negligible. We denote the corresponding excess number of $e$-folds as $\Delta N_{e}$.
@@
For $\alpha/f=60$, inflation ends when the potential and kinetic energies of the axion field become comparable and the backreaction of the gauge field becomes important post-inflation. The $\alpha/f=30$ and $50$ cases have the similar behavior. However, for $\alpha/f=75$ and $90$, the amplification of the gauge fields is significant to backreact on the evolution of the axion field during inflation itself. When the backreaction of gauge fields on the axion field dynamics becomes significant during inflation, it reduces the axion field velocity, thereby extending the duration of inflation. The duration of inflation extends by $\Delta N_{e}=1.7$ $e$-folds for $\alpha/f=75$ and $\Delta N_{e}=4.0$ $e$-folds for $\alpha/f=90$ compared to cases with smaller coupling strengths. Our results are similar to those obtained in ref. [53], except that here the extended duration of inflation $\Delta N_{e}$ is slightly shorter.
@@
In scenarios where the backreaction of gauge fields is negligible during inflation, it ends when the potential and kinetic energy become comparable. However, when gauge field backreaction is significant during inflation, it ends when the gauge field energy density is comparable to the axion potential energy.
@@
Figure 1: The time evolution of different energy densities (top row), $\xi$ (middle row), and the first slow-roll parameter, $\epsilon_{H}$ (bottom row), is shown for $\alpha/f=60$ (left column), 75 (middle column), and 90 (right column). The vertical gray lines indicate the end of inflation for each case. In the middle and bottom rows, the violet lines represent the time evolution of $\xi$ and $\epsilon_{H}$ for the homogeneous $\phi$ case. In panels (e) and (f), the dot-dashed gray lines show the analytical GEF results discussed in refs. [45, 51].
@@ 3.3 Comparison with previous work in the homogeneous $\phi$ case
Figure 2: The spectra of the magnetic and electric components of the gauge field at different times are shown for $\alpha/f=75$ (left, blue curves, run D) and $\alpha/f=90$ (right, red curves, run E).
@@
As seen in figure 2, both magnetic and electric energy spectra initially peak at a scale close to the minimum wave number of the simulation box. As time progresses, the spectral peaks shift towards higher wave numbers, eventually reaching close to the maximum box wave number for run E. To extend the simulation to later times, a larger simulation box would be required.

### v2

# Lattice simulations of axion-U(1) inflation: gravitational waves, magnetic fields, and scalar statistics

@@ Abstract
We numerically study axion-U(1) inflation, focusing on the regime where the coupling between axions and gauge fields results in significant backreaction from the amplified gauge fields during inflation. These amplified gauge fields not only generate high-frequency gravitational waves (GWs), but also enhance spatial inhomogeneities in the axion field. GWs serve as key probe for constraining the coupling strength between the axion and gauge fields. We find that, when backreaction is important during inflation, the constraints on the coupling strength due to GW overproduction are relaxed compared to previous studies, in which backreaction matters only after inflation. Moreover, our results suggest that the probability density function (PDF) of axion fluctuations tends toward a Gaussian distribution even in cases where gauge field backreaction is important only after inflation. This aligns with previous studies where the same effect was observed for cases with strong backreaction during inflation. This finding can be crucial for future studies of primordial black hole (PBH) formation, which can further constrain the coupling strength. We also calculate the spectrum of the produced magnetic fields in this model and find that their strength is compatible with the observed lower limits.
@@ 1 Introduction
The idea of inflation originated from attempts to solve the singularity problem [1] and to address the horizon, isotropy and flatness problems in the hot big bang along with the monopole problem of grand unification, see [2, 3, 4] and [5, 6]. Inflation not only ameliorates the latter set of issues, but most importantly provides a natural mechanism [7, 8, 9] for generating initial density perturbations in the universe from mere vacuum fluctuations, see also [10, 11, 12, 13]. These perturbations account for anisotropies in the cosmic microwave background radiation (CMB) recently measured by Planck collaboration [14] with unprecedented precision, thereby confirming the predictions of this mechanism and the theory of inflation. Later on these perturbations collapse to form all galaxies and the large-scale structure in the universe. Following the initial proposal, numerous models have been suggested to realize an inflationary era in the early universe (for details see refs. [15, 16, 17, 18, 19]). Axion inflation models are particularly interesting because they provide a natural high energy physics ingredient for a long lasting inflation - an approximate shift symmetry of the potential [20]. After the first proposal, several models have been suggested in the context of axion inflation [21, 22, 23, 24, 25, 26, 27, 28, 29, 30]; see ref. [31] for a review. One of these models [25] considers a coupling of the axion field with the U(1) gauge field and we focus on this model in this paper; see ref. [32] for the observational status of axion-U(1)inflation. The axion inflation model, where the axion couples to a gauge field, exhibits a rich phenomenology. This includes the production of gravitational waves (GWs) [33, 34, 35, 36, 37, 38, 39, 40], primordial black holes (PBHs) [41, 42, 43], matter-antimatter asymmetry [44, 45], and primordial magnetic fields [46, 47, 48, 49, 50, 51, 52, 53]. Axion inflation with SU(2) gauge fields and its phenomenology has also been extensively explored [28, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71]. The rich phenomenology of axion inflation has prompted extensive studies to understand the dynamics of these models. In scenarios with a sufficiently strong coupling between the axion and the gauge field, the backreaction of the produced gauge fields becomes important either during inflation or post inflation depending on the coupling strength. This model including the backreaction of the gauge fields has been studied semi-analytically through perturbative approaches [72, 73, 74] and the gradient expansion formalism (GEF) [75, 53] under the assumption of a homogeneous axion field. More recently, efforts have been made to numerically investigate the strong backreaction regime, where the assumption of homogeneous axion field is relaxed [76, 77, 78]. Motivated by the importance of inhomogeneities of the axion field, as demonstrated in the numerical studies, these inhomogeneities have also been incorporated perturbatively into GEF [79].
@@
This rich phenomenology of axion inflation has been used to constrain the coupling strength between the axion and gauge fields. The coupling between the axion and the gauge field results in the production of gauge fields, which in turn generate high-frequency GWs, typically in the GHz range [38, 39]. These GWs contribute to the early universe’s total energy budget as additional radiation degrees of freedom. The bound on additional radiation degrees of freedom from CMB observations constrains the produced GWs, which further limits the coupling between the axion and the gauge fields. We reexamine the bounds on the coupling strength between the axion and the gauge field, considering stronger couplings than those explored in previous works [38, 39]. In [38], the authors considered scenarios where the backreaction due to the produced gauge field became important in the post-inflationary era but not during inflation. For one of their models, the chaotic inflation model, they found that the coupling strength should be below $(m_{\mathrm{pl}}/70)^{-1}$ for an inflaton mass equal [^1] Here and throughout the paper $m_{\mathrm{pl}}=1.22\times 10^{19}\,\,\rm{GeV}$ represents the Planck mass. to $1.2\times 10^{-6}m_{\mathrm{pl}}$. In our study, we explore scenarios with larger coupling strengths, where the gauge field energy density becomes large enough to affect the duration of inflation. In these scenarios, inflation ends when the gauge field energy density becomes comparable to the inflaton’s potential energy. Earlier studies of such scenarios, presented in refs. [76, 77], did not address the production of GWs that we investigate in this paper. Our findings indicate that in such scenarios, the bounds on the coupling strength are relaxed compared to previous estimates [38, 39].
@@
Additionally, the produced gauge field can also enhance spatial fluctuations of the axion field. If these spatial fluctuations exceed a threshold value, they can lead to the formation of PBHs222 For other models of the formation of PBHs during inflation see, e.g., refs. [80, 81, 82, 83, 84]. in the post-inflationary era. This phenomenon has been analyzed in detail considering the $\chi^{2}$-distribution of the spatial fluctuations of the axion field. Given that the backreaction of the gauge field on the axion field is quadratic in nature, and assuming the gauge field has a Gaussian distribution, the spatial fluctuations of the axion field are expected to follow a $\chi^{2}$-distribution. By using constraints on PBH abundances, a bound on the axion and gauge field coupling has been obtained [41, 42]. A recent numerical study of axion inflation [76] considers the case when the backreaction of the gauge field on the axion evolution becomes significant during inflation. It was found that the distribution of the spatial fluctuations of the axion field starts to deviate from the $\chi^{2}$-distribution and begins to resemble a Gaussian distribution. In this paper, we study the probability distribution of spatial fluctuations in the axion field over a range of coupling strengths. We find that the distribution tends to a $\chi^{2}$-distribution during the evolution. However, it gradually transitions to a Gaussian distribution at a later time, irrespective of whether the backreaction of the gauge fields becomes important during or after inflation.
@@ 2 Axion-U(1) inflation model
Here, $\phi$ represents the axion field and $V(\phi)$ denotes its potential, which in our analysis is taken as [^3] Although chaotic inflation has been ruled out by CMB observations [18], we have chosen this model for several reasons. First, it simplifies our analysis. Second, it allows us to compare our results directly with those of previous studies. Lastly, many other inflationary potentials can be approximated by a quadratic potential near the end of inflation, and our study primarily focuses on this late phase of the inflationary period. $V(\phi)=m^{2}\phi^{2}/2$ with $m=1.06\times 10^{-6}m_{\rm pl}$, $F_{\mu\nu}=\partial_{\mu}A_{\nu}-\partial_{\nu}A_{\mu}$ represents the gauge field strength tensor, [^4] In this work, we identify the gauge fields with the Standard Model hypercharge sector. The production of fermions is important in this context [85, 53], but their numerical modeling touches upon many unsolved questions, which we have to leave for future work. Alternatively, the scenario in this study could apply to gauge fields in a dark sector, but in that case, these dark sector gauge fields would not provide magnetogenesis as discussed in section 5. while $\tilde{F}^{\mu\nu}$ is its Hodge dual, $\alpha/f$ denotes the coupling strength between the axion field and the U(1) gauge fields, and $R$ is the Ricci scalar. Using this, and denoting derivatives with respect to conformal time $\eta$ by a prime, we get the following dynamical equations [^5] In this paper, we neglect the role of metric perturbations. The terms that couple metric perturbations to the axion field are suppressed by slow-roll parameters and can thus be ignored during inflation [86, 87]. However, the role of these terms may become important near the end of inflation, which we leave for future investigation; see refs. [88, 89, 78] for recent studies on this topic. for the axion-gauge field evolution [49]
@@ 3.1 Initial conditions
In this section, we describe the initial conditions for the axion and gauge fields used in our simulations. The initial conditions are set well within the inflationary era. To achieve accelerated expansion of the universe, the pressure $p$ must satisfy $p<-\rho/3$ which, in the absence of gauge fields, implies $\phi^{\prime 2}<a^{2}V(\phi)$. Using this condition in equations (2.5), we get
@@
where $H=\dot{a}/a=\mathcal{H}/a$ is the usual Hubble parameter, dot denotes derivative with respect to cosmological time $dt=ad\eta$ and $\epsilon_{H}$ represents the first slow-roll parameter. The above expressions imply
@@ 3.2 Dynamics of axion and gauge fields: inferred from simulations
To study the dynamics of the axion-U(1) inflation model, we perform a series of simulations with different coupling strengths [^7] It is to be noted that our coupling strengths are in the units of the inverse of the Planck mass. However, many previous studies use the inverse of the reduced Planck mass for normalization. To convert our coupling strengths to the reduced Planck mass normalization, our values need to be divided by $\sqrt{8\pi}$. table 1 provides the values of the coupling strengths in both units for clarity. $\alpha/f=30$, $50$, $65$, $75$, and $90$. The parameters for these simulations are given in table 1. In figure 1, we present the evolution of different energy densities for $\alpha/f=60$, $75$, and $90$. We show the evolution of the kinetic energy density, $\rho_{\mathrm{kin}}$ (dashed blue curves), potential energy density, $V(\phi)$ (solid cyan curves), and gauge field energy density, $\rho_{g}$ (dot-dashed orange curves) in panels (a), (b), and (c) of figure 1 for the coupling strengths $\alpha/f=60$, $75$, and $90$, respectively. Panels (d), (e), and (f) illustrate the evolution of $\xi$, while panels (g), (h), and (i) display the first slow-roll parameter, $\epsilon_{H}$, for the corresponding cases. Inflation ends when $\epsilon_{H}=1$. The vertical gray lines indicate the epoch of the end of inflation in each case. In this figure, $e$-folds $N$ is defined [^8] As usual we define dimensionless $e$-folds $N$ through the relation $dN=Hdt$ with cosmological time $t$ and Hubble parameter $H$. in such a way that $N=0$ would be the end of inflation had one neglected backreaction, even in scenarios where backreaction is significant. We denote the corresponding excess number of $e$-folds as $\Delta N_{e}$.
@@
For $\alpha/f=60$, inflation ends when the potential and kinetic energies of the axion field become comparable and the backreaction of the gauge field becomes important post-inflation. The $\alpha/f=30$ and $50$ cases have a similar behavior. However, for $\alpha/f=75$ and $\alpha/f=90$, the amplification of the gauge fields is significant to backreact on the evolution of the axion field during inflation itself. When the backreaction of gauge fields on the axion field dynamics becomes significant during inflation, it reduces the axion field velocity, thereby extending the duration of inflation. The duration of inflation extends by $\Delta N_{e}=1.7$ $e$-folds for $\alpha/f=75$ and $\Delta N_{e}=4.0$ $e$-folds for $\alpha/f=90$ compared to cases with smaller coupling strengths. Our results are similar to those obtained in ref. [77], except that here the extended duration of inflation $\Delta N_{e}$ is slightly shorter.
@@
In scenarios where the backreaction of gauge fields is negligible during inflation, it ends when the potential and kinetic energy become comparable. However, when gauge field backreaction is significant during inflation, it ends when the gauge field energy density is comparable to the axion potential energy. This conclusion is based on the numerical results of the simulations shown in figure 1. Similar findings have also been reported in the simulations of ref. [77].
@@
Figure 1: The time evolution of different energy densities (top row), $\xi$ (middle row), and the first slow-roll parameter, $\epsilon_{H}$ (bottom row), is shown for $\alpha/f=60$ (left column), $\alpha/f=75$ (middle column), and $\alpha/f=90$ (right column). The vertical gray lines indicate the end of inflation for each case. In the middle and bottom rows, the violet lines represent the time evolution of $\xi$ and $\epsilon_{H}$ for the homogeneous $\phi$ case. In panels (e) and (f), the dot-dashed gray lines show the analytical GEF results discussed in refs. [51, 75].
@@ 3.3 Comparison with previous work in the homogeneous $\phi$ case
Figure 2: The spectra of the magnetic and electric components of the gauge field at different times are shown for $\alpha/f=75$ (left, blue curves, run D) and $\alpha/f=90$ (right, red curves, run E). Here, the cross (‘x’) represents the wavenumber corresponding to the Hubble horizon at the time of the spectrum.
@@
As seen in figure 2, both magnetic and electric energy spectra initially peak at a scale close to the minimum wave number of the simulation box. As time progresses, the spectral peaks shift towards higher wave numbers, eventually reaching close to the maximum box wave number for run E. To extend the simulation to later times, a larger simulation box would be required. For the simulation shown in figure 2, we note that there is a late-time growth in the electric and magnetic energy spectra for superhorizon wave numbers. A similar growth has also been seen in other simulations [77, 91].